# A Bias-Free Training Paradigm for More General AI-generated Image Detection

### v1

@@ 10 Calibration Metrics
11 Additional Ablation
In this Section we further investigate the influence of our training dataset. First of all, we consider a SoTA method, RINE [21], that is designed to exploit also low-level forensic features and compare different strategies. The model is trained on three different datasets: ProGAN [43], Latent Diffusion (LDM) [10] and the dataset with content augmentation we proposed in this work (inpainted+). Results are presented in Tab. 7 and show that our dataset is particularly helpful to obtain better and more calibrated results. On average we achieve an improvement of around 7% in terms of AUC and balanced accuracy compared to training on LDM, and a much larger gain, around 20% in terms of accuracy, with respect to ProGAN. The advantage seems to be significant especially for synthetic generators that are much different from Latent Diffusion, such as DALL-E 3, Firefly and FLUX.
@@ 11 Additional Ablation
Ours 94.7 / 88.3 93.6 / 88.1 90.3 / 80.5 95.3 / 89.3 89.4 / 79.1 95.4 / 88.1 91.6 / 83.2 97.1 / 91.5 95.4 / 88.5 96.3 / 90.1 93.9 / 86.7
Table 9: Performance on each generator included in FakeBench dataset in terms of AUC and balanced Accuracy.
We also conduct and additional experiment and train our network, SigLip [49], on the very recent Diffusion-generated Deepfake Detection (D3) dataset [3], of about 8M synthetic images (from 256 $\times$ 256 to 1024 $\times$ 1024) from the generators SD 1.4, SD 2.1, SDXL, and DeepFloyd IF. The images are generated using prompts taken from the description of the real source from LAION (text-driven generation). We can notice that although the AUC is similar ($\simeq$ 95%), there is a significant increase in terms of balanced accuracy (70.7 vs. 89.9) and decrease in terms of both NLL (1.21 vs 0.28) and ECE (0.28 vs 0.07). This confirms that our training paradigm enables better calibration and improved generalization.
12 Additional Generalization Analysis
@@
Evaluation on same generators from different datasets. Here we further expand on the analysis conducted in Fig. 2 of the main paper where we have shown that some detectors achieve different performance on the same generator when the images are taken from two different datasets. This puzzling behavior suggests the possibility that these methods rely on subtle dataset biases besides true traces left by the synthetic generator. In Fig. 10 we extend this analysis to all SoTA methods described in Sec. 9. More specifically, we analyze the performance in terms of AUC and balanced accuracy over three synthetic generators: Midjourney, SDXL and DALL-E 3 that come from three different datasets PolarDiffShield [23], Synthbuster [2] and FakeInversion [7]. As said before, for several methods the performance is not consistent on the same generator and can vary even by 20% from one dataset to another. In addition, for some methods the AUC is around 50%, which corresponds to random choice, or even below 50% which means that the detector tends to invert the labels between real and fake.
@@
13 Additional Robustness Analysis
Here we want to further investigate the performance of our method compared to SOTA in terms of robustness. To this end, we analyze two variants of our method, content augmentation and diverse category (inpainted+) and with additional standard augmentation operations (inpainted++).
Analysis of post-processing operations. We show further results on FakeInversion [7], where the real images have been retrieved from the web. To better understand the effect of compression and resizing we compare the performance when applying such operations. In particular, to simulate the upload on social networks, we resize with a scale factor randomly sampled between 0.7 and 1, and compress with a JPEG quality factor between 70 and 100. In Tab. 10 results show that the performance on such dataset drops substantially, except for DMID and our method, though our approach inpainted++ has slightly better calibration metrics.
Analysis on content shared on-line. Fig. 11 illustrates the balanced accuracy on real/fake images shared on the web, with multiple versions of the same image (as already shown for the proposal in Fig. 7 of the main paper). Accuracy is evaluated over a 100-day period from the time of initial publication of each image, with times on a logarithmic scale. We compare our proposal with the best performing SoTA methods on this dataset of around 1400 images. We can notice that the performance drops after only one day, after which most competitors are stuck below 65%. Only the variant inpainted++ of the proposed method that comprises more aggressive augmentation is able to ensure an average accuracy around 85% even after many days from the first on-line post.
Figure 11: Results of SoTA detectors on real and fake images that went viral on internet. The performance is in term of balanced accuracy evaluated from the initial online post (Log scale).
@@ 13 Additional Robustness Analysis
DMID 97.3 / 96.1 0.25 / .041 94.3 / 81.3 0.55 / .182
LGrad 84.3 / 77.2 2.27 / .200 60.1 / 55.4 7.59 / .426
UnivFD 54.9 / 52.8 2.19 / .391 49.7 / 49.9 2.57 / .434
DeFake 69.8 / 63.3 0.95 / .225 69.3 / 62.6 0.98 / .236
DIRE 53.3 / 51.8 13.4 / .360 55.4 / 51.6 13.4 / .358
AntifakePrompt     -    / 53.9     -    /     -     -    / 54.7     -    /     -
NPR 91.5 / 87.0 4.96 / .123 43.3 / 49.9 27.1 / .501
FatFormer 68.1 / 59.7 3.45 / .386 48.6 / 50.3 5.44 / .490
FasterThanLies 49.7 / 48.6 3.64 / .476 50.9 / 49.9 3.13 / .458
RINE 69.6 / 63.6 4.84 / .319 62.3 / 52.9 6.28 / .437
AIDE 85.5 / 76.9 0.54 / .137 67.2 / 56.3 0.93 / .276
LaDeDa 91.6 / 84.7 3.03 / .129 51.9 / 53.1 24.8 / .454
C2P-CLIP 74.1 / 59.6 0.82 / .260 71.8 / 59.0 0.89 / .284
CoDE 87.5 / 78.7 0.74 / .143 82.5 / 74.4 0.89 / .171
Ours (inpainted+) 94.0 / 79.6 0.76 / .182 89.3 / 77.3 0.80 / .176
Ours (inpainted++) 95.8 / 82.0 0.44 / .154 93.7 / 82.5 0.42 / .122

### v2

@@ Appendix C Calibration Metrics
Appendix D Additional Ablation
In this Section we further investigate the influence of our training dataset. We train our network on the very recent Diffusion-generated Deepfake Detection (D3) dataset [3], of about 8M synthetic images (from 256 $\times$ 256 to 1024 $\times$ 1024) from the generators SD 1.4, SD 2.1, SDXL, and DeepFloyd IF. The images are generated using prompts taken from the description of the real source from LAION (text-driven generation). From Tab. 8, we can notice that although the AUC is similar, there is a significant increase in terms of balanced accuracy (87.5 vs 95.2) and decrease in terms of both NLL (0.31 vs 0.14) and ECE (0.13 vs 0.04). This confirms that our training paradigm enables better calibration and improved generalization.
@@ Appendix D Additional Ablation
B-Free (ours) 99.3 / 96.4 97.7 / 88.5 99.3 / 95.2 96.5 / 86.5 95.5 / 87.7 100. / 98.9 98.7 / 94.9 99.9 / 98.7 100. / 98.7 98.0 / 94.5 98.5 / 94.0
Table 10: Performance on each generator included in FakeBench dataset in terms of AUC and balanced Accuracy. Bold underlines the best performance for each column with a margin of 1%.
Appendix E Additional Generalization Analysis
@@
Evaluation on same generators from different datasets. Here we further expand on the analysis conducted in Fig. 2, where we have shown that some detectors achieve different performance on the same generator when the images are taken from two different datasets. This puzzling behavior suggests the possibility that these methods rely on subtle dataset biases besides true traces left by the synthetic generator. In Fig. 10 we extend this analysis to all SoTA methods described in Sec. B. More specifically, we analyze the performance in terms of AUC and balanced accuracy over three synthetic generators: Midjourney, SDXL and DALL-E 3 that come from three different datasets PolarDiffShield [25], Synthbuster [2] and FakeInversion [7]. As said before, for several methods the performance is not consistent on the same generator and can vary even by 20% from one dataset to another. In addition, for some methods the AUC is around 50%, which corresponds to random choice, or even below 50% which means that the detector tends to invert the labels between real and fake.
@@
Finally, we generated a set of 2,000 images with two autoregressive models [24, 51]. whose architecture is significantly different from the one used to generate the training data. On both models we obtain good results with an average of 99% in terms of AUC and 94.2% in terms of Accuracy. This is probably due to the similarity between the tokenizer used in these models and the latent embedders of Stable Diffusion models.
Figure 11: Robustness analysis in terms of balanced Accuracy carried out on nine generators of Synthbuster under three different post-processing operations: JPEG compression, resizing and blurring. The five best performing SoTA detectors on Synthbuster have been included in the analysis.
Appendix F Additional Robustness Analysis
We repeat the experiment reported in Figure 6 to test robustness under various operations and carry out comparisons with SOTA methods. Results are shown in Fig. 11 under three post-processing operations: JPEG compression, resizing, and blurring. We can observe that our approach is more robust by a large margin compared with other methods and can ensure a balanced accuracy that is always above 80% even in the most challenging scenario.
Finally we further investigate robustness performance on FakeInversion [7], where the real images have been retrieved from the web. To better understand the effect of compression and resizing we compare the performance when applying such operations. In particular, to simulate the upload on social networks, we resize with a scale factor randomly sampled between 0.7 and 1, and compress with a JPEG quality factor between 70 and 100. In Tab. 11 results show that the performance on such datase drops substantially, except for DMID and our method, though our approach has better calibration metrics.
@@ Appendix F Additional Robustness Analysis
DMID 97.3 / 96.1 0.25 / .041 94.4 / 81.3 0.55 / .182
LGrad 84.4 / 77.2 2.27 / .200 60.2 / 55.4 7.59 / .426
UnivFD 54.9 / 52.8 2.19 / .391 49.7 / 50.0 2.57 / .434
DeFake 69.8 / 63.3 0.95 / .225 69.3 / 62.7 0.98 / .236
DIRE 53.3 / 51.8 13.4 / .360 55.5 / 51.7 13.4 / .358
AntifakePrompt    -    / 53.9    -    /    -    -    / 54.8    -    /    -
NPR 91.6 / 87.0 4.96 / .123 43.3 / 49.9 27.1 / .501
FatFormer 68.2 / 59.7 3.45 / .386 48.7 / 50.4 5.44 / .490
FasterThanLies 49.7 / 48.6 3.64 / .476 51.0 / 49.9 3.13 / .458
RINE 69.6 / 63.6 4.84 / .319 62.4 / 52.9 6.28 / .437
AIDE 85.5 / 76.9 0.54 / .137 67.3 / 56.4 0.93 / .276
LaDeDa 91.7 / 84.7 3.03 / .129 51.9 / 53.1 24.8 / .454
C2P-CLIP 74.1 / 59.6 0.82 / .260 71.9 / 59.0 0.89 / .284
CoDE 87.5 / 78.7 0.74 / .143 82.5 / 74.4 0.89 / .173
B-Free (ours) 99.3 / 86.3 0.32 / .144 98.5 / 86.4 0.32 / .131